OpenStudy (anonymous):

Prove by induction N E i = (N(N+1))/2 i = 0

OpenStudy (anonymous):

Have you done proofs via induction before?

OpenStudy (anonymous):

\[\sum_{i=0}^{N} i = \frac{ N(N+1) }{ 2 }\] That should be more clear.

OpenStudy (anonymous):

Yeah, got it.

OpenStudy (anonymous):

So have you done proofs via induction before? Because they can seem pretty weird before you get used to them...

OpenStudy (anonymous):

A little bit, I've seen a couple youtube videos, but those are not like this one.

OpenStudy (anonymous):

Have you ever done Logic before?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well, there's one type of argument form called "modus ponens" (you don't have to know that) that helped me learn Induction. It goes like this: If P, then Q P Therefore, Q

OpenStudy (anonymous):

Oh, yes then. Sorry.

OpenStudy (anonymous):

That looks familiar is what I mean

OpenStudy (anonymous):

So, as an example, I might say something like: If it's raining, then I'll go to bed. It's raining. And with those two statements, you can conclude that: I'm going to bed. You with me?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Well, that's one way you can look at Induction (and, I'm mentioning it because it was how things started to click with me regarding Induction).

OpenStudy (anonymous):

So it goes something like this: "Let's just assume for a minute that the sum of the first "k" positive whole numbers is equal to k(k+1)/2." That's all you're doing, just assuming that it is true. Then, that would mean that the sum of the first k+1 numbers is equal to: \[= \frac{ k(k+1) }{ 2 } + (k+1)\]

OpenStudy (anonymous):

With me?

OpenStudy (anonymous):

Yes, I think so

OpenStudy (anonymous):

Now, just factor out the "k+1": \[= (k+1)(\frac{ k }{ 2 }+1)\] Which can be simplified further as: \[= \frac{ (k+1)(k+2) }{ 2 }\]

OpenStudy (anonymous):

And that's the exact same form of the equation we started with--the only difference is that that is for the sum of the first "k+1" positive integers and not "k". That's the "If P, Then Q" part I was talking about--you show that if you assume it's true for the first "k" numbers, then it also must be true for "k+1" numbers. Once you've gotten that, the rest is easy. You just show that it's true when k=1, which is obvious because the sum of the first "1" positive whole numbers is, of course, 1, and when you plug in "1" into your initial formula you get: \[\frac{ 1(1+1) }{2 }= 1\] And that's your "P" part.

OpenStudy (anonymous):

Thus, you've proven two steps: 1.) If it's true for n=k, then it's true for n=k+1 2.) it's true for n=1. And all at once you've proven that that formula is true for EVERY SINGLE POSITIVE INTEGER.

OpenStudy (anonymous):

Ok, that helped! thank you :)

OpenStudy (anonymous):

Ok, hope that makes sense--don't feel bad if it's still a little foggy (you wouldn't believe how long it took me to get a hang of Induction...). Good luck!